water was added to 100g of a salt mgcl2 to produce 80.0cm3 of saturated solution at 25°c. its solubility at 25°c is 10 mol/dm 3. calculate the mass of the salt which remains undissolved
1
Expert's answer
2020-10-01T07:53:06-0400

Solubility refers to the mass of a solute per 100 g of a solvent .

can be express as g/100 g of water.

converting mass of salt into mole.

1 mole of cacl2 = 40 + 2(35.5) = 121g


"100g * (1mol\/121g) = 0.826 mol"


solution of into dm3

"80cm^3 *(1dm^3\/1000cm^3) = 0.08dm^3"


solubility = 0.826mol /0.08dm3 =10.325 mol/dm3


subtracting the actual solubility with that of the salt added

at the 25oc 10.325 mol/dm3 - 10.0mol/dm3 = 0.325mol


0.325 mol is equivalent to

"0.325mol* (121g\/1mol) = 39.325 g" this is the mass that remained un dissolved